# Approximating the coefficients of the Bessel functions

### v4

@@ 6. Leading order terms for the type A Dunkl bilinear form
$n_{\nu^{\prime}}(x)\triangleq\sum_{\begin{subarray}{c}\pi\in NC(|\nu^{\prime}|),\,\gamma(\pi)=\nu^{\prime}\end{subarray}}W_{\pi}^{A}(x).$
@@
$\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}),\\ \gamma(\pi)=\gamma((\nu_{j}:j\in S_{i}))\end{subarray}}W_{\pi}^{A}(N\theta)$
$=N^{\ell(\lambda)}\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\prod_{i=1}^{\ell(\lambda)}\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}),\\ \gamma(\pi)=\gamma_{i}\end{subarray}}W_{\pi}^{A}(N\theta)$
$=N^{\ell(\lambda)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}W_{\pi}^{A}(N\theta).$
@@ 7. Leading order terms for the type BC Dunkl bilinear form
7.1. Proof of part (B) of Theorem˜1.1
@@
7.2. Leading order terms of the type BC Dunkl bilinear form in the $\theta_{0}N\rightarrow c_{0}\in\mathbb{C}$, $\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regime
@@
$W^{BC}(\pi)(x,y)\triangleq\prod_{i\in[k],\,b(i;\,\pi)=0}(x+\mathbf{1}\{d(i;\,\pi)\text{ is odd}\}y+d(i;\,\pi)).$
@@
$[p_{\lambda},p_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=$ $N^{\ell(\lambda)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}W^{BC}(\pi)(2N\theta_{0},2\theta_{1})\prod_{B\in\pi}x_{|B|}$
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{BC}_{\nu\lambda}(2c_{0},2c_{1})[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right).$
@@
See the proof of Theorem˜6.17. For the implication of (a) from (b), consider the formula for $W^{BC}([l])(2c_{0},2c_{1})$ for some $l\in 2\mathbb{N}$. If $d(i;\,\pi)$ is odd, then the term $2c_{0}+2c_{1}+d(i;\,\pi)$ is nonzero. Moreover, if $d(i;\,\pi)$ is even and at least two, then the term $2c_{0}+d(i;\,\pi)$ is nonzero; note that if $d(i;\,\pi)=0$, then we must have that $b(i;\,\pi)=1$. Therefore, $W^{BC}([l])(2c_{0},2c_{1})\not=0$, so $\mathcal{W}^{BC}(2c_{0},2c_{1})[k]$ is invertible for all $k\geq 1$ by Lemma˜7.7. ∎
Corollary 7.9 ([rectangularmatrix]*Theorems 4.8 and 5.5).
@@
    $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}W^{BC}(\pi)(2c_{0},2c_{1})\prod_{B\in\pi}|B|c_{(|B|)}.$
@@ 8.1. Orthogonality results
Assume that $\theta\in\mathbb{C}$ and $\mathcal{D}(D^{N}(\theta))$ is invertible. Suppose $a\in\mathbb{C}^{N}$. The coefficients of terms of $J_{a}^{D^{N}(\theta)}(x_{1},\ldots,x_{N})$ with all odd degrees in $x_{1},\ldots,x_{N}$ are all zero if and only if $a_{1}\cdots a_{N}=0$.
@@
The following result shows that $e$ is a harmonic function with respect to $\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))$ for all $\theta_{0},\theta_{1}\in\mathbb{C}$.
Lemma 8.8.
For all $\theta_{0},\theta_{1}\in\mathbb{C}$, $\mathcal{D}(BC^{N}(\theta_{0},\theta_{1}))(p_{(2)})e=0$.
Proof.
First, observe that
$\mathcal{D}_{1}e=(1+2\theta_{1})x_{2}\cdots x_{N}.$
Afterwards, it is straightforward to deduce that $\mathcal{D}_{1}^{2}e=0$. ∎
Corollary 8.9.
For all $\theta\in\mathbb{C}$, $\mathcal{D}(D^{N}(\theta))(p_{(2)})e=0$.
Proof.
Set $\theta_{1}=0$ in Lemma˜8.8. ∎
8.2. The leading order terms of $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$ and an introduction to the leading order terms of $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ for $\lambda,\nu\in\Gamma_{\text{even}}$
Next, we prove the results analogous to those of Section˜6 for the type D root system. However, we must now compute both $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$ and $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$ for $\lambda,\nu\in\Gamma$. First, we consider the computation of $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$. As we will see, these values are asymptotically equivalent to the values of $p_{\lambda},p_{\nu}]_{A^{N-1}(2\theta)}$ in both the $|\theta N|\rightarrow\infty$ and $\theta N\rightarrow c\in\mathbb{C}$ regimes.
First, we consider the $|\theta N|\rightarrow\infty$ regime.
Theorem 8.10.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then, Theorem˜6.1 with $A^{N-1}(\theta)$ replaced by $D^{N}(\theta)$ and $\theta$ replaced by $2\theta$ is true.
Proof.
We can use the same method as the proof of Theorem˜6.1. The reason for this is that because the switches are to distinct indices that are not elements of $\{j_{i}:1\leq i\leq\ell(\lambda)\}$ to achieve the leading order term, the switch from $j_{i}$ to $j$ is equivalent to $2\theta d_{j_{i}}$. This is because after applying $\frac{1-\sigma_{j_{i}j}}{x_{j_{i}}-x_{j}}$ or $\frac{1-\tau_{j_{i}}\tau_{j}s_{j_{i}j}}{x_{j_{i}}+x_{j}}$ to $x_{j_{i}}^{d}$ for some $d\geq 1$, the only resulting term without $x_{j}$ is $\theta x_{j_{i}}^{d-1}$ in both cases. Hence, we only need to replace $\theta$ with $2\theta$ in Theorem˜6.1 to compute the leading order term of $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$. ∎
Define the infinite dimensional matrix $\mathcal{M}^{D}\in\mathbb{Z}_{\geq 0}^{\Gamma_{\text{even}}\times\Gamma_{\text{even}}}$ such that for $\lambda,\nu\in\Gamma_{\text{even}}$, $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$.
Next, we consider the $\theta N\rightarrow c\in\mathbb{C}$ regime.
Theorem 8.11.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}|\theta N|=\infty$. Then, statements (a) and (c) of Theorem˜8.21 are equivalent.
Proof.
This follows from Theorem˜8.10 and the proof of Theorem˜6.10 after $\Gamma$ and $\mathcal{M}^{A}$ are replaced with $\Gamma_{\text{even}}$ and $\mathcal{M}^{D}$, respectively. ∎
Theorem 8.12.
Suppose $k\geq 1$, $\lambda,\nu\in\Gamma_{\text{even}}[k]$, and $\ell(\lambda)\leq\ell(\nu)$. Then, Theorem˜6.14 with $A^{N-1}(\theta)$ replaced by $D^{N}(\theta)$ and $N\theta$ replaced by $2N\theta$ is true.
Proof.
Similarly to the proof of Theorem˜8.10, since the switches are to distinct indices that are not elements of $\{j_{i}:1\leq i\leq\ell(\lambda)\}$ to achieve the leading order term, the switch from $j_{i}$ to $j$ is equivalent to $2\theta d_{j_{i}}$. Then, to compute the leading order term of $p_{\lambda},p_{\nu}]_{D^{N}(\theta)}$, we only need to replace $N\theta$ with $2N\theta$ in Theorem˜6.14. ∎
Theorem 8.13.
Suppose $F_{N}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}c_{\lambda}(N)p_{\lambda}\right)+e\exp\\ \left(\sum_{\lambda\in\Gamma_{\text{even};\,N}}d_{\lambda}(N)p_{\lambda}\right)$. Assume that $\lim_{N\rightarrow\infty}\theta N=c\in\mathbb{C}$. Then, in Theorem˜8.26, (a) implies (c) and if $2c$ is not a negative integer, then (c) implies (a).
Proof.
This follows from Theorem˜8.12 and the proof of Theorem˜6.17 after $\Gamma$ and $c$ replaced with $\Gamma_{\text{even}}$ and $2c$, respectively. ∎
Remark 8.14.
Theorems 8.10, 8.11, 8.12, and 8.13 also follow from the analogous results for the type BC root system after setting $\theta_{1}=0$.
Next, we consider the computation of $ep_{\lambda},ep_{\nu}]_{D^{N}(\theta)}$, which is considerably more complicated. First, we exhibit a more direct computation in the case where $p_{\lambda},p_{\nu}=p_{(2)}^{k}$ and the multiplicity function is nonnegative by using Theorem˜8.1.
Lemma 8.15.
Suppose $k\geq 0$. Then, for all $\theta\in\mathbb{R}_{\geq 0}$,
$\left[ep_{(2)}^{k},ep_{(2)}^{k}\right]_{D^{N}(\theta)}=\frac{\int_{\mathbb{R}^{N}}\prod_{i=1}^{N}|x_{i}|^{2}\prod_{j=i+1}^{N}|x_{i}+x_{j}|^{2\theta}|x_{i}-x_{j}|^{2\theta}wdx}{\int_{\mathbb{R}^{N}}\prod_{i=1}^{N-2}\prod_{j=i+1}^{N}|x_{i}+x_{j}|^{2\theta}|x_{i}-x_{j}|^{2\theta}wdx}\left[p_{(2)}^{k},p_{(2)}^{k}\right]_{BC^{N}(\theta,1)}.$
Proof.
Suppose $\theta\geq 0$. By Theorem˜8.1,
$\left[ep_{(2)}^{k},ep_{(2)}^{k}\right]_{D^{N}(\theta)}=c_{N}^{-1}\int_{\mathbb{R}^{N}}(\mathcal{D}(e^{-p_{(2)}})ep_{(2)}^{k})(\mathcal{D}(e^{-p_{(2)}})ep_{(2)}^{k})h^{2}wdx,$
where $c_{N}$, $h$, and $w$ are defined with respect to the type $D$ root system with constant multiplicity $\theta$.
By Lemma˜8.8, $e$ is harmonic. Then, using [dunkl_integralkernel]*Proposition 3.9 gives that
$\mathcal{D}(e^{-p_{(2)}})ep_{(2)}^{k}=(-1)^{k}k!2^{k}L_{k}^{\left(\frac{3N}{2}+\gamma-1\right)}\left(\frac{p_{(2)}}{2}\right)e,$
where $\gamma=N(N-1)\theta$ and $L_{n}^{(\alpha)}$ denotes the generalized Laguerre polynomial. Hence,
$\left[ep_{(2)}^{k},ep_{(2)}^{k}\right]_{D^{N}(\theta)}=c_{N}^{-1}\int_{\mathbb{R}^{N}}\left(k!2^{k}L_{k}^{\left(\frac{3N}{2}+\gamma-1\right)}\left(\frac{p_{(2)}}{2}\right)\right)^{2}(eh)^{2}wdx.$
For a general positive root system $\mathcal{R}^{+}$ and multiplicity function $\theta$, $\gamma\triangleq\sum_{r\in\mathcal{R}^{+}}\theta(r)$. Let $\gamma^{\prime}$, $h^{\prime}$, and $c_{N}^{\prime}$ denote the values of $\gamma$, $h$, and $c_{N}$, respectively, with respect to $B^{N}(\theta_{0},\theta_{1})$; note that we specify that we are using the type B root system, although using the type C root system is equivalent. We similarly have that
$\left[p_{(2)}^{k},p_{(2)}^{k}\right]_{BC^{N}(\theta,1)}=(c^{\prime}_{N})^{-1}\int_{\mathbb{R}^{N}}\left(k!2^{k}L_{k}^{\left(\frac{N}{2}+\gamma^{\prime}-1\right)}\left(\frac{p_{2}}{2}\right)\right)^{2}(h^{\prime})^{2}wdx$
However, it is evident that $\gamma^{\prime}=\gamma+N$ and $(h^{\prime})^{2}=(eh)^{2}$, if we assume that the positive roots for the type B root system are $\{[\mathbf{1}\{i=j\}]_{j\in[N]}^{T}:i\in[N]\}$ in addition to the positive roots for the type D root system. Then, we have that
$\left[ep_{(2)}^{k},ep_{(2)}^{k}\right]_{D^{N}(\theta)}$ $=\frac{c^{\prime}_{N}}{c_{N}}\left[p_{(2)}^{k},p_{(2)}^{k}\right]_{BC^{N}(\theta,1)}$
$=\frac{\int_{\mathbb{R}^{N}}\prod_{i=1}^{N-1}|x_{i}|^{2}\prod_{j=i+1}^{N}|x_{i}+x_{j}|^{2\theta}|x_{i}-x_{j}|^{2\theta}wdx}{\int_{\mathbb{R}^{N}}\prod_{i=1}^{N}\prod_{j=i+1}^{N}|x_{i}+x_{j}|^{2\theta}|x_{i}-x_{j}|^{2\theta}wdx}\left[p_{(2)}^{k},p_{(2)}^{k}\right]_{BC^{N}(\theta,1)}.$

### v5

@@ 6.3. Leading order terms of the type A Dunkl bilinear form in the high temperature regime
$n_{\nu^{\prime}}(x)\triangleq\sum_{\begin{subarray}{c}\pi\in NC(|\nu^{\prime}|),\,\gamma(\pi)=\nu^{\prime}\end{subarray}}W^{A}(\pi)(x).$
@@
$\prod_{i=1}^{\ell(\lambda)}\pi((\nu_{j}:j\in S_{i}))\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}),\\ \gamma(\pi)=\gamma((\nu_{j}:j\in S_{i}))\end{subarray}}W^{A}(\pi)(N\theta)$
$=N^{\ell(\lambda)}\sum_{\nu=\gamma_{1}+\cdots+\gamma_{\ell(\lambda)}}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\prod_{i=1}^{\ell(\lambda)}\sum_{\begin{subarray}{c}\pi\in NC(\lambda_{i}),\\ \gamma(\pi)=\gamma_{i}\end{subarray}}W^{A}(\pi)(N\theta)$
$=N^{\ell(\lambda)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}\prod_{B\in\pi}x_{|B|}W^{A}(\pi)(N\theta).$
@@ 7. Leading order terms for the type BC Dunkl bilinear form
7.1. Proof of part (B) of Theorem˜1.2
@@
7.2. Leading order terms of the type BC Dunkl bilinear form in the high temperature regime
@@
$W^{BC}(\pi)(x,y)\triangleq\prod_{i\in[k],\,b(i;\,\pi)=0}(2x+\mathbf{1}\{d(i;\,\pi)\text{ is odd}\}2y+d(i;\,\pi)).$
@@
$[p_{\lambda},p_{\nu}]_{BC^{N}(\theta_{0},\theta_{1})}=$ $N^{\ell(\lambda)}\prod_{l=1}^{\ell(\nu)}\nu_{l}\pi(\nu)\left[\prod_{l=1}^{\ell(\nu)}x_{\nu_{l}}\right]\prod_{i=1}^{\ell(\lambda)}\sum_{\pi\in NC(\lambda_{i})}W^{BC}(\pi)(N\theta_{0},\theta_{1})\prod_{B\in\pi}x_{|B|}$
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N^{\ell(\nu)}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=\sum_{\lambda\in\Gamma_{\text{even}}}\mathcal{W}^{BC}_{\nu\lambda}(c_{0},c_{1})[p_{\lambda}]\exp\left(\sum_{\gamma\in\Gamma_{\text{even}}}c_{\gamma}p_{\gamma}\right).$
@@
See the proof of Theorem˜6.17. For the implication of (a) from (b), consider the formula for $W^{BC}([l])(c_{0},c_{1})$ for some $l\in 2\mathbb{N}$. If $d(i;\,\pi)$ is odd, then the term $2c_{0}+2c_{1}+d(i;\,\pi)$ is nonzero. Moreover, if $d(i;\,\pi)$ is even and at least two, then the term $2c_{0}+d(i;\,\pi)$ is nonzero; note that if $d(i;\,\pi)=0$, then we must have that $b(i;\,\pi)=1$. Therefore, $W^{BC}([l])(c_{0},c_{1})\not=0$, so $\mathcal{W}^{BC}(c_{0},c_{1})[k]$ is invertible for all $k\geq 1$ by Lemma˜7.7. ∎
We similarly have that Theorem˜7.8 generalizes the results of the paper [rectangularmatrix] while Corollary˜7.9 has already been proved in the paper.
Corollary 7.9.
@@
    $\lim_{N\rightarrow\infty}\frac{1}{(\theta N)^{|\nu|}}[p_{\nu},F_{N}]_{BC^{N}(\theta_{0},\theta_{1})}=\prod_{i=1}^{\ell(\nu)}\sum_{\pi\in NC^{\text{even}}(\nu_{i})}W^{BC}(\pi)(c_{0},c_{1})\prod_{B\in\pi}|B|c_{(|B|)}.$
@@ 8.1. Orthogonality results
Assume that $\theta\in\mathbb{C}$ and $\mathcal{D}(D^{N}(\theta))$ is invertible. Suppose $a\in\mathbb{C}^{N}$. The coefficients of the terms of $J_{a}^{D^{N}(\theta)}(x_{1},\ldots,x_{N})$ with all odd degrees in $x_{1},\ldots,x_{N}$ are all zero if and only if $a_{1}\cdots a_{N}=0$.
@@
8.2. Proof of part (C) of Theorem˜1.2